HIGH RESOLUTION ROVIBRATIONAL SPECTROSCOPY OF JET-COOLED PHENYL RADICAL: THE ν_{19} OUT-OF-PHASE SYMMETRIC C-H STRETCH

\underline{GRANT T. BUCKINGHAM}, CHIH-HSUAN CHANG, and DAVID J. NESBITT, JILA, National Institute of Standards and Technology and University of Colorado, Department of Chemistry and Biochemistry, Boulder, CO 80309.

Phenyl radical has been studied via sub-Doppler infrared spectroscopy in a slit supersonic discharge expansion source, with assignments for the highest frequency b_{2} out-of-phase C-H symmetric stretch vibration (ν_{19}) unambiguously confirmed by ≤ 6 MHz (0.0002 cm^{−1}) agreement with microwave ground state combination differences of McMahon et al. [Astrophys. J. 590, L61-64 (2003)]. Least squares analysis of > 100 resolved rovibrational peaks in the sub-Doppler spectrum to a Watson Hamiltonian yields precision exited-state rotational constants and a vibrational band origin (ν_{0} = 3071.8915(4) cm^{−1}) consistent with a surprisingly small red-shift (0.9 cm^{−1}) with respect to Ar matrix isolation studies of Ellison and coworkers [J. Am. Chem. Soc. 123, 1977 (2001)]. Nuclear spin weights and inertial defects confirm the vibrationally averaged planarity and ^{2}A_{1} rovibronic symmetry of phenyl radical, with analysis of the rotational constants consistent with a modest C_{2v} distortion of the carbon backbone frame due to partial sp rehybridization of the σ C radical-center. Most importantly, despite the number of atoms (N = 11) and vibrational modes (3N−6 = 27), phenyl radical exhibits a remarkably clean jet cooled high resolution IR spectrum that shows no evidence of intramolecular vibrational relaxation (IVR) phenomena such as local or non-local perturbations due to strongly coupled nearby dark states. This provides strong support for the feasibility of high resolution infrared spectroscopy in other cyclic aromatic hydrocarbon radical systems.